OpenStudy (samigupta8):

The number of integral order pair (x,y) that satisfy the system of equation |x+y-4|=5 and |x-3|+|y-1|=5 is /are

OpenStudy (samigupta8):

@hartnn

OpenStudy (anonymous):

why don't you use elimination or substitution method

OpenStudy (samigupta8):

How many times??

OpenStudy (samigupta8):

For that i have to repeatedly put up the values and see whether they satisfy or not...

OpenStudy (welshfella):

yea thats a bit tedious

OpenStudy (samigupta8):

Actually!! It is...that's why posted it..

OpenStudy (anonymous):

ok i would give you the answer if am wrong tell me

OpenStudy (anonymous):

(x+y−4)=5

OpenStudy (anonymous):

x-intercept: (9,0) y-intercept: (0,9)

OpenStudy (samigupta8):

It is mod here...

OpenStudy (samigupta8):

| | symbol is modulus

OpenStudy (anonymous):

that means every answer is positive

OpenStudy (samigupta8):

No!

OpenStudy (samigupta8):

X ,y can be negative too!

OpenStudy (samigupta8):

@ganeshie8

OpenStudy (samigupta8):

@baru

OpenStudy (samigupta8):

@Michele_laino

OpenStudy (freckles):

the only way I can think of to do this without looking at a calculator is to... think of all the possible cases: \[\text{ Case 1: } x+y-4 \ge0 \text{ ; } x-3\ge0 \text{ ; }y-1 \ge 0\] There are 7 more cases: The first is assume +,+,+ (which I have above) Second is +,+,- Third is +,-,+ Fourth is +,-,- Fifth is -,+,- Sixth is -,+,+ Seventh is -,-,+ Eighth is -,-,- I will show you the rest of the first case... So assuming the first case we have the following equations: \[x+y-4=5 \\ x-3+y-1=5 \\ \text{ simplifying } \\ x+y=9 \\ x+y=9 \\ \text{ so we only have to look at } x+y=9\] \[\text{ now remember in this case we have } x \ge 3 \text{ and }y \ge 1\] So x+y can be the following: 3+6 4+5 5+4 6+3 7+2 8+1 I got no further because y would be less than 1 and notice I started with x is 3... since that is the least number x can be anyways... if you were to continue you would go to Case 2: \[\text{ Case 2 : }=x+y-4 \ge0 \text{ ; } x-3 \ge 0 \text{ ; } y-1 \le 0 \\ \text{ so this would make your equations look like } \\ x+y-4=5 \text{ and } x-3-(y-1)=5 \\ x+y=9 \text{ and } x-y-2=5 \\ x+y=9 \text{ and } x-y=7 \\ \\ \text{ solve the system } \\ 2x=16 \\ x=8 \\ y=1 \\ \text{ make sure this fits in with our assumptions; it does }\] anyways continue with the other cases... I'm trying to think if there is a way to rule out some of these cases... Like that last case gave us a x and y we already had...

OpenStudy (samigupta8):

Well what i think is we can see the similarity between the two mod terms . Both are 5 ...so we can write the first term as |x-3+y-1|=5

OpenStudy (samigupta8):

It was |x-3+y-1|=5 And second term is |x-3|+|y-1|=5 So that means the first term is simply |x-3+y-1|=|x-3|+|y-1|

OpenStudy (ikram002p):

|x+y-4|=5 and |x-3|+|y-1|=5 |x+y-4|=5 |x-3+y-1|=5 assume x>3,y>1 |x-3|+|y-1|=5 has (4,5),(5,4),(6,3),(7,2),(8,1),(3,6)

OpenStudy (ikram002p):

be careful ... this is only the first set of solutions i think.. i like what @freckles did but a bit of too much..

OpenStudy (samigupta8):

We can say that this cN be true only when either x-3>=0 and y-1>=0 Or x-3<=0 and y-1<=0

OpenStudy (samigupta8):

I m talking about my method ...

OpenStudy (freckles):

well I think @samigupta8 is right on her equation if you use triangle inequality \[5=|x-3+y-1| \le |x-3|+|y-1|=5 \\ \text{ but this inequality only makes sense if } \\ |x-3+y-1|=|x-3|+|y-1|\]

OpenStudy (freckles):

so everything she has I think can be reduce to solving the equation \[|x+y-4|=|x-3|+|y-1|\] but also make sure after solving for x and y that |x+y-4|=5

OpenStudy (samigupta8):

Sure!!

OpenStudy (freckles):

@ikram002p is your way more reduced than my way?

OpenStudy (freckles):

like less cases?

OpenStudy (freckles):

8 cases is so much :p

OpenStudy (ikram002p):

same boat they are :P...

OpenStudy (freckles):

even after applying that triangle inequality I see no way out of getting out of those 8 cases

OpenStudy (samigupta8):

We have simplified our equation like this X-3>=0 and y-1>=0 Which gives 6 ordered pairs ...

OpenStudy (samigupta8):

And a case when both are less than 0

OpenStudy (freckles):

I think I'm going to look at just this for now solving for integer u and v such that the following are true: \[|u+v|=5 \\ |u|+|v|=5\] I think this will be easier to look at

OpenStudy (samigupta8):

I got 12 cases...

OpenStudy (freckles):

Squaring first equation: \[(u+v)^2=25 \\ u^2+v^2+2uv=25 \] Squaring second equation: \[(|u|+|v|)^2=25 \\ u^2+v^2+2|u||v|=25\] \[2uv=2|u||v| \\ uv=|u||v| \\ uv=|uv| \text{ is true only if } uv \ge 0 \\ u=x-3 \text{ while } v=y-1 \\ \text{ so we need \to solve } (x-3)(y-1) \ge 0\] so x-3 is positive and y-1 is positive or x-3 is negative and y-1 is negative but they can't have different sign So I think we can reduce this down to 2 cases maybe if I haven't made some error

OpenStudy (samigupta8):

Correct!! Hereby you will get 6 cases for u and v both positive And 6 cases for u and v both negative

OpenStudy (freckles):

oh answers is what you meant yes 12 answers looks right

OpenStudy (samigupta8):

Ryt ..